# Single-Shot Metric Depth from Focused Plenoptic Cameras

### v1

@@ Abstract
Our work explores the potential of focused plenoptic cameras for dense metric depth. We propose a novel pipeline that predicts metric depth from a single plenoptic camera shot by first generating a sparse metric point cloud using machine learning, which is then used to scale and align a dense relative depth map regressed by a foundation depth model, resulting in dense metric depth. To validate it, we curated the Light Field & Stereo Image Dataset [^1] Available at zenodo.org/records/14224205. (LFS) of real-world light field images with stereo depth labels, filling a current gap in existing resources. Experimental results show that our pipeline produces accurate metric depth predictions, laying a solid groundwork for future research in this field.
@@ I Introduction
From a single monocular camera, however, we can only estimate an up-to-scale 3D reconstruction. The true scale of a scene is unobservable from multi-view geometric models [3]. In learning-based single-view approaches, it also is unobservable for self-supervised models [4], as their losses are based on the same multi-view geometric models. For supervised ones, the scale accuracy depends on the training and test data [5]. Multi-sensor setups, such as visual-inertial, stereo, or multi-camera, do observe the scale. However, in addition to their higher degree of hardware complexity, they require a precise alignment and calibration, and their accuracy is limited by the baseline between the cameras, which is limited by the applications. In the case of visual-inertial systems, their state may not be observable for certain motions.
Light field cameras represent a promising alternative for metric depth estimation. Their inner configuration involves a microlens array (MLA) placed between the main lens and the sensor. This allows the acquisition not only of the intensity but also the direction of light rays that interact with the sensor, by acquiring range-dependent sub-aperture images arranged in a grid pattern [6]. This unique internal structure allows the capture of multiple views and ranges of local areas with a single device and a single shot, overcoming the limitations of traditional monocular and multi-sensor setups. Additional advantages include their higher light yield (as they produce focused images even with large lens apertures), the elimination of disocclusion issues (since the scene is captured in 3D using a single main lens), and the absence of moving parts (unlike optical refocusing systems). These features have already found applications in robotics, such as on-orbit servicing and robotic exploration [7, 8].
@@
In this paper, we explore the potential of a focused plenoptic camera for single-view metric depth estimation (see Fig. 1). Specifically, we present an end-to-end pipeline that leverages light field images to resolve the scale ambiguity of a learning-based monocular depth regressor, hence producing dense metric depth maps. We believe that our approach, which leverages both, priors from a foundation model with global receptive fields and geometric cues between multiple microlenses, is bound to outperform geometric triangulation methods from local correspondences in plenoptic images. To benchmark our approach, we curated a novel real-world Light Field & Stereo Image Dataset (LFS) that we will release upon acceptance. Our results show that we outperform related baselines, specifically the manufacturer’s software for depth, sparse and presumably based only on geometry, and the state-of-the-art foundation model Depth Anything.
@@ II Related Work
Single-view depth is geometrically ill-posed [19]. Various computational methods, inspired by human visual perception, have been developed to tackle this challenge using visual cues such as perspective and and visual appearance through lighting and occlusion [20, 21]. Deep neural networks have shown impressive results at predicting per-pixel depth or disparity by learning image patterns from databases with and without supervision [22, 4, 23, 24, 5, 25].
@@
Depth from light field cameras. Capturing light field images in a single exposure combines the simplicity of single-shot setups with the depth information coming from local calibrated multiple views, thereby potentially adding scale information to single-view methods. They have evolved from the plenoptic camera 1.0, which focused the main lens at the microlens array (MLA) plane, allowing for post-capture processing but at a lower resolution, to the plenoptic camera 2.0, which focuses the MLA on the main lens’s image plane, enhancing image quality by balancing spatial and angular resolution. For both devices, however, a recurring challenge is the scarcity of available data. This is due to the limited availability of light field cameras and the difficulties in setting up ground-truth measuring systems for supervised learning. While camera arrays offer an effective answer, they are costly and impractical for widespread use. Additionally, plenoptic cameras are hindered by their reliance on proprietary software, high costs, and limited datasets [26, 27, 28]. Some works have addressed the depth estimation task from light field capturing devices [29, 30, 31, 32]. However, when exploring learning-based approaches, a recurring obstacle emerges: the scarcity of available data. This shortage stems from the limited availability of light field capturing devices and from the difficulty of obtaining precise ground truth, as it requires an external system and data registration. Moreover, the technology incorporated into these cameras is conditioned by the dependence on the manufacturer’s proprietary software for decoding raw data [28]. In addition, both their low-volume production and their manufacturing complexity contribute to their elevated cost, preventing widespread adoption [27].
@@
Existing light field datasets, such as the Stanford Light Field Archives [33, 34, 35], provide valuable resources but often lack alignment with the more modern plenoptic 2.0 cameras, limiting their usability. Although other datasets with similar setups exist [36, 37], their utility is constrained by insufficient data volume, and differences in data formats and hardware version. In addition, the lack of real-world datasets for plenoptic 2.0 cameras has led most research to rely on synthetic datasets, which, while valuable, lack the realism needed for robust deployment [38, 39, 40, 41, 42, 43].
@@
Figure 4: Overview of our single-shot depth pipeline from light field images. The Image Processing Toolkit pre-processes captured data, which is then used to train the Microlens Depth Network for sparse metric depth estimation. These are used to refine the inference by Depth Anything, producing a dense metric depth map. Filtered stereo depth serves as the ground truth, following densification and scale alignment.
@@ III The Light Field & Stereo (LFS) Dataset
Hardware and software. The cameras are mounted on a robust mechanical framework made of aluminium profiles. The setup includes two Allied Vision Mako G-419C cameras [44] and a Raytrix R5 plenoptic camera [45], which is based on the Baumer HXG40c model. The Mako cameras are compact, industrial vision cameras, while the Raytrix R5 is a light field camera with up to 1 MP effective output resolution and 4.2 Megarays light field resolution. All cameras feature a CMOS CMV4000 sensor at 4.2 MP resolution, supporting 25 fps under the GigE vision standard. The three cameras are connected to a notebook through a 2.5G switch and a 2.5G Ethernet adapter. To ensure overlapping fields of view between the Mako stereo pair and the plenoptic camera, the cameras are positioned close together with rigid alignment and matching lenses. Synchronization is achieved by using the plenoptic camera’s exposure signal as a master trigger for the stereo. The configuration, control, and data processing utilizes the vendors’ SDKs, which rely on GenICam transport layer (GenTL) libraries to communicate with the cameras. This setup is extended with in-house packages for concurrent deployment of processes, frame distribution via shared memory, image post-processing, and efficient storage. Geometric camera calibration is carried out using the method by Zhang, Sturm, and Maybank [46, 47] implemented in the camera calibration toolbox DLR CalLab [48], along with the stepwise plenoptic camera calibration method [49] and the extrinsic calibration method by Strobl [50]. The RxLive software from Raytrix is also utilized to capture plenoptic images and visualize scene representations, with custom methods developed for pre-processing plenoptic images. Metric stereo depth is estimated by an in-house implementation of the SGM algorithm [51]. Lastly, custom scripts were developed for stereo image reprojection.
@@
Plenoptic depth. The information obtained from RxLive is processed as an alternative ground-truth depth source for benchmarking purposes. This involves decoding the virtual depth data values obtained from the manufacturer, transforming them into metric space using a parametrized thin lens camera model [49].
@@
TABLE I: Comparison of our Light Field Sparse and Dense depths against baselines, A Random Depth generator, the Sparse Depth provided by the Raytrix manufacturer software, and the foundation deep model Depth Anything. Note how both our sparse and dense modules outperform the chosen baselines.
@@ IV Single-Shot Depth from Plenoptic Cameras
We present a complete pipeline to obtain dense, metric depth maps from a single light field image. The optical configuration of the plenoptic camera is leveraged to treat the microlens array as a multi-view stereo, providing the disparity to estimate metric depth. The proposed pipeline combines image processing, microlens-scale depth estimation (Section IV-A), and subsequent densification and refinement to generate a dense metric depth map (Section IV-B), see Fig. 4.
@@ IV-A Sparse Depth from Plenoptic Images
Our model follows an encoder-decoder convolutional architecture. Flower stacks are provided to the encoder in the form of 4D tensors $X_{i}$ of size $(N,C,H,W)$, for which $N$ is the batch size, $C$ the number of channels, and $H\!\times\!W$ the height and width of each microlens projection. The proposed model infers a single metric depth value per stack, corresponding to the prediction at the centroid of the main microlens. This network, which we denote as Microlens Depth Network, is composed of 2D-convolution layers, batch normalization, activation functions, and fully-connected bottleneck layers. It captures spatial relationships among microlenses, leveraging the redundancy among the microimages in the flower stack to enhance depth estimation robustness, e.g. in regions with occlusions. The estimated depths are then used to create a sparse depth map aligned with the central sub-aperture image, i.e., at a lower resolution due to the limited number of microlenses (8,837) with respect to the sensor’s resolution.
@@ IV-B Dense Depth from Plenoptic Cameras
The depth map obtained as described in Section IV-A is sparse due to the limited number of microlenses, is inaccurate in regions with weak texture, which we also remove from the depth map. In order to obtain dense per-pixel metric depth values, we fuse these results with the pretrained model Depth Anything [25], which generates dense disparity maps from monocular images. Depth Anything is based on the Vision Transformer (ViT) architecture and trained on extensive datasets, ensuring robustness across a variety of scenarios, and showing state-of-the-art performance in the most recent benchmarks [52]. However, Depth Anything’s disparity maps are not in metric scale. We use a robust regression method to align them to the sparse metric depth values predicted by our Microlens Depth Network. We use the Theil-Sen estimator [53, 54] for this alignment, as it handles outliers in an effective manner. First, dense disparity values are extracted at locations with corresponding sparse metric depths (the latter are converted to disparities for consistency, to achieve a monotonic function that preserves the order), and then aligned using the Theil-Sen approach. This non-parametric method calculates the median slope $m$ from all pairs of points $(x_{i},y_{i})$ and $(x_{j},y_{j})$ as follows:
@@ IV-C Implementation Details
Loss function. We use the pixel-wise mean squared error (MSE) between our depth predictions $\hat{y}$ and their ground-truth values $y$
@@
Figure 6: Light Field Dense Depth results. From top to bottom: the natural image of a scene (first) allows Depth Anything to densely (yet up-to-scale) infer depth (third). Then, scale alignment using metric values from our Sparse Light Field Depth transforms it into a dense metric depth map of the scene (fourth). Note the closer alignment in the color scale of our Light Field Dense depth with the ground truth depth (second) than that of Depth Anything.
@@ V-B Training Details
We train the Microlens Depth Network for 125 epochs using batches of 128 flower stacks. This requires around 10 hours using a Quadro FV100 Volta GPU. We optimize using Adam[55] with a learning rate of $0.001$ and apply early stopping with 20 epochs patience. For Depth Anything, we used $1024\!\times\!1024$ RGB images. Theil-Sen regression was applied to align scale and offset using 11,002 valid depth values out of 17,321 sparse data points. The process took 1.78 seconds per image on the same GPU.
@@ V-C Comparison against Baselines
Our Microlens Depth Network inference took 10.54 minutes for a test set of 84,650 flower stacks, representing 20 % of the LFS dataset (10 images). The test images encompass diverse scenes with elements at varying distances. Table I shows the aggregated results, along with those of related baselines.
@@
TABLE II: Error metrics for additional methods explored. The first and fourth rows present the proposed methods. Each of these is then ablated and compared quantitatively. The proposed outperform the alternatives.
@@ VI Conclusion
This work introduces a novel approach for single-shot depth from a plenoptic camera 2.0. We developed a novel Microlens Depth Network that successfully infers sparse metric depth from plenoptic images, we then synthesize a natural image from the plenoptic one, that we forward pass through the foundation single-view up-to-scale dense depth model Depth Anything. Finally, we regress the true scale and offset values to correct the dense output from our sparse depth. Our end-to-end pipeline demonstrates the feasibility of generating dense metric depth maps from single shots. Our work makes several key contributions, including the design of a specialized neural network for depth estimation, the creation and release of the Light Field & Stereo Image Dataset (LFS) – a new dataset with plenoptic images and corresponding reprojected metric depth labels – and the development of a comprehensive image pre-processing methodology suited for learning-based applications. These contributions advance the state of the art in light field imaging and single-view depth estimation, establishing a foundation for further research.

### v2

@@ Abstract
Our work explores the potential of focused plenoptic cameras for dense metric depth. We propose a novel pipeline that predicts metric depth from a single plenoptic camera shot by first generating a sparse metric point cloud using a neural network, which is then used to scale and align a dense relative depth map regressed by a foundation depth model, resulting in a dense metric depth. To validate it, we curated the Light Field & Stereo Image Dataset [^1] Dataset available at https://zenodo.org/records/14224205. (LFS) of real-world light field images with stereo depth labels, filling a current gap in existing resources. Experimental results show that our pipeline produces accurate metric depth predictions, laying a solid groundwork for future research in this field.222 Work partially supported by the DLR Impulse Project SaiNSOR.
@@ I Introduction
From a single pinhole camera, we can estimate the 3D reconstruction of an unknown scene using monocular structure-from-motion (SfM), but this is subject to scale ambiguity [3]. In learning-based single-view approaches, scale is also unobservable for self-supervised models [4], as their losses are based on the same multi-view geometric constraints. For supervised ones, the scale accuracy depends on the training and test data [5]. Multi-sensor setups, such as visual-inertial, stereo, or multi-camera, do observe the scale. However, in addition to their higher degree of hardware complexity, they require a precise alignment and calibration, and their accuracy is limited by the baseline between the cameras, which is conditioned by the applications. In the case of visual-inertial systems, their state may not be observable for certain motions.
Light field cameras represent a promising alternative for metric depth estimation. Their inner configuration involves a microlens array (MLA) placed between the main lens and the sensor. This allows the acquisition not only of the intensity but also the direction of light rays that interact with the sensor, by acquiring range-dependent sub-aperture images arranged in a grid pattern [6]. This unique internal structure allows the capture of multiple views and ranges of local areas with a single device and a single shot, overcoming the limitations of traditional monocular and multi-sensor setups. Additional advantages include their higher light capture (as they produce focused images even with large lens apertures), the elimination of disocclusion issues (since the scene is captured in 3D using a single main lens), and the absence of moving parts (unlike optical refocusing systems). These features have already found applications in robotics, such as on-orbit servicing and robotic exploration [7, 8].
@@
In this paper, we explore the potential of a focused plenoptic camera for single-view metric depth estimation (see Fig. 1). Specifically, we present an end-to-end pipeline that leverages light field images to resolve the scale ambiguity of a learning-based monocular depth regressor, hence producing dense metric depth maps. We believe that our approach, which leverages both, priors from a foundation model with global receptive fields and geometric cues between multiple microlenses, is bound to outperform geometric triangulation methods from local correspondences in plenoptic images. To benchmark our approach, we curated a novel real-world Light Field & Stereo Image Dataset (LFS) that we release with this paper. Our results show that we outperform related baselines, specifically the manufacturer’s software for depth, sparse and presumably based only on geometry, and the state-of-the-art foundation model Depth Anything.
@@ II Related Work
Single-view depth is geometrically ill-posed [19]. Various computational methods, inspired by human visual perception, have been developed to tackle this challenge using visual cues such as perspective and and visual appearance through lighting and occlusion [20, 21]. Deep neural networks have shown impressive results at predicting per-pixel depth or disparity by learning image patterns from databases with and without supervision [22, 4, 23, 24, 25, 5, 26].
@@
Depth from light field cameras. Capturing light field images in a single exposure combines the simplicity of single-shot setups with the depth information coming from local calibrated multiple views, thereby potentially adding scale information to single-view methods. They have evolved from the plenoptic camera 1.0, which focused the main lens at the microlens array (MLA) plane, allowing for post-capture processing but at a lower resolution, to the plenoptic camera 2.0, which focuses the MLA on the main lens’s image plane, enhancing image quality by balancing spatial and angular resolution. For both devices, however, a recurring challenge is the scarcity of available data. This is due to the limited availability of light field cameras and the difficulties in setting up ground-truth measuring systems for supervised learning. While camera arrays offer an effective answer, they are costly and impractical for widespread use. Additionally, plenoptic cameras are hindered by their reliance on proprietary software, high costs, and limited datasets [27, 28, 29].
Some works have addressed the depth estimation task using classical methods developed for earlier light field capturing devices [30, 31, 32, 33]. However, there are no available depth estimation methods for the plenoptic camera 2.0, primarily due to the lack of a publicly available geometric model and the scarcity of available real-world data, which impedes the development of learning-based approaches. This shortage stems from the limited availability of light field capturing devices and from the difficulty of obtaining precise ground truth, as it requires an external system and data registration. Moreover, the technology incorporated into these cameras is conditioned by the dependence on the manufacturer’s proprietary camera model and software for decoding raw data [29]. In addition, both their low-volume production and their manufacturing complexity contribute to their elevated cost, preventing widespread adoption [28].
@@
Existing light field datasets, such as the Stanford Light Field Archives [34, 35, 36], provide valuable resources but often lack alignment with modern plenoptic 2.0 cameras, limiting their usability. Although other datasets with similar setups exist [37, 38], their utility is constrained by insufficient data volume, different data formats and hardware version. Hence, the lack of real-world datasets for plenoptic 2.0 cameras has led most research to rely on synthetic datasets, which, while valuable, lack the realism needed for robust deployment [39, 40, 41, 42, 43, 44].
Depth completion methods generate dense depth maps from sparse representations, e.g. in the context of LIDAR data [45]. These methods typically refine and densify sparse depth maps using unguided techniques, such as interpolation or hand-crafted features, within the same modality [46, 47, 48]. However, image-guided methods have proven more successful, especially with noisy depth data [49, 50, 51]. In the case of a light field camera, sparse depth maps are generated from a limited number of microlenses in the sensor. Integrating foundation models as an analogous image-guided method offers a promising approach for depth completion in our domain [52].
@@
Figure 4: Overview. The Image Processing Toolkit pre-processes captured data, which is then used to train the Microlens Depth Network for sparse metric depth estimation. These are used to refine the inference by Depth Anything, producing a dense metric depth map. Filtered stereo depth serves as the ground truth, following densification and scale alignment.
@@ III The Light Field & Stereo (LFS) Dataset
Hardware and software. The cameras are mounted on a robust mechanical framework made of aluminium profiles. The setup includes two Allied Vision Mako G-419C cameras [53] and a Raytrix R5 plenoptic camera [54], which is based on the Baumer HXG40c model. The Mako cameras are compact, industrial vision cameras, while the Raytrix R5 is a light field camera with up to 1 MP effective output resolution and 4.2 Megarays light field resolution. All cameras feature a CMOS CMV4000 sensor at 4.2 MP resolution, supporting 25 fps under the GigE vision standard. The three cameras are connected to a notebook through a 2.5G switch and a 2.5G Ethernet adapter. To ensure overlapping fields of view between the stereo pair and the plenoptic camera, they are positioned close together with rigid alignment and matching lenses. Synchronization is achieved by using the plenoptic camera’s exposure signal as a master trigger for the stereo. The configuration, control, and data processing utilizes the vendors’ SDKs, which rely on GenICam transport layer (GenTL) libraries to communicate with the cameras. This setup is extended with in-house packages for concurrent deployment of processes, frame distribution via shared memory, image post-processing, and efficient storage. Geometric camera calibration is carried out using the method by Zhang, Sturm, and Maybank [55, 56] implemented in the camera calibration toolbox DLR CalLab [57], along with the stepwise plenoptic camera calibration method [58] and the extrinsic calibration method by Strobl and Hirzinger [59]. The RxLive software from Raytrix is also utilized to capture plenoptic images and visualize scene representations, with custom methods developed for pre-processing plenoptic images. Metric stereo depth is estimated by an in-house implementation of the SGM algorithm [60]. Lastly, custom scripts were developed for stereo image reprojection.
@@
Plenoptic depth. The data obtained from RxLive is processed as an alternative ground-truth depth source for benchmarking purposes. This involves decoding the virtual depth values obtained from the manufacturer and transforming them into metric ones using a thin lens camera model [58].
@@
TABLE I: Comparison of our Light Field Sparse and Dense depth against 1) A Random Depth generator, 2) the Sparse Depth provided by the Raytrix manufacturer software, and 3) the foundation model Depth Anything.
@@
Light Field Sparse Depth (our Microlens Depth Network) 124.68 5.55 52.75 2.63 84.83 88.40 91.25 0.3949
@@ IV Single-Shot Depth from Plenoptic Cameras
We present a complete pipeline to obtain dense, metric depth maps from a single light field image, that combines image processing, microlens-scale depth estimation (Section IV-A), and subsequent densification and refinement to generate a dense metric depth map (Section IV-B), see Fig. 4.
@@ IV-A Sparse Depth from Plenoptic Images
Our model follows an encoder-decoder convolutional architecture. Flower stacks are provided to the encoder in the form of 4D tensors $X_{i}$ of size $(N,C,H,W)$, for which $N$ is the batch size, $C$ the number of channels, and $H\!\times\!W$ the height and width of each microlens projection onto the sensor. Our model infers a single metric depth value per stack, corresponding to the prediction at the centroid of the main microlens of each stack. This network, which we denote as Microlens Depth Network, is composed of 2D convolutions, batch normalization, activation functions, and fully-connected bottleneck layers. It captures spatial relationships among microlenses, leveraging their redundancy in the flower stack to enhance depth estimation robustness, e.g. in regions with occlusions. The estimated depths form a sparse depth map aligned with the central sub-aperture image at a lower resolution due to the limited number of microlenses (8,837). The resulting depth map is inaccurate in regions with weak textures, which we then remove by applying a texture filter.
@@ IV-B Dense Depth from Plenoptic Cameras
In order to obtain dense per-pixel metric depth values, we fuse these results with the pretrained model Depth Anything [26], which generates dense disparity maps from monocular images. Depth Anything is trained on extensive datasets, ensuring robustness across a variety of scenarios, and showing state-of-the-art performance in the most recent benchmarks [61]. However, its disparity maps are not in metric scale. We use a robust regression method to align them to the sparse metric depth values predicted by our Microlens Depth Network. We use the Theil-Sen estimator [62, 63] for this alignment, as it handles outliers in an effective manner. First, dense disparity values are extracted at locations with corresponding sparse metric depths (the latter are converted to disparities for consistency, to achieve a monotonic function that preserves the order), and then aligned using the Theil-Sen approach. This non-parametric method calculates the median slope $m$ from all pairs of points $(x_{i},y_{i})$ and $(x_{j},y_{j})$ as follows:
@@ IV-C Implementation Details
Loss. We use the mean squared error (MSE) between our per-pixel depth predictions $\hat{y}$ and their ground-truth values $y$
@@
Figure 6: Qualitative results. From natural images (first row), Depth Anything infers dense depth (yet up-to-scale) (third row). Then, scale alignment with metric values from our Sparse Light Field Depth transforms it into a dense metric depth (fourth row). Note our Light Field Dense depth aligns more closely with the ground truth depth (second row) than Depth Anything.
@@ V-B Training Details
We train the Microlens Depth Network for 125 epochs using batches of 128 flower stacks. This requires around 10 hours using a Quadro FV100 Volta GPU. We optimize using Adam [64] with a learning rate of $0.001$. For Depth Anything, we used $1024\!\times\!1024$ RGB images. Theil-Sen regression was applied to align scale and offset using 11,002 valid depth values out of 17,321 sparse data points. The process took 1.78 seconds per image on the same GPU.
@@ V-C Comparison against Baselines
Our Microlens Depth Network took 10.54 minutes for a test set of 84,650 flower stacks, representing 20% of the LFS dataset (10 images). The test images encompass diverse scenes with elements at varying distances. Table I shows the aggregated results, along with those of related baselines.
@@
TABLE II: Error metrics for additional methods explored. The first and fourth rows present the proposed methods. Each of these is then ablated and compared quantitatively. The proposed ones outperform the alternatives.
@@ VI Conclusion
This work introduces a novel approach for single-shot depth from a plenoptic camera 2.0. We developed a novel Microlens Depth Network that successfully infers sparse metric depth from plenoptic images, we then synthesize a natural image from the plenoptic one, that we forward pass through the foundation single-view up-to-scale dense depth model Depth Anything. Finally, we regress the true scale and offset values to correct the dense output from our sparse depth. Our end-to-end pipeline demonstrates the feasibility of generating dense metric depth maps from single shots. Our work makes several key contributions, including the design of a specialized neural network for depth estimation, the creation and release of the Light Field & Stereo Image Dataset (LFS) – a new dataset with plenoptic images and corresponding reprojected metric depth labels – and the development of a comprehensive image pre-processing methodology suited for learning-based applications. These contributions advance the state of the art in light field imaging and single-view depth estimation, establishing a foundation for further research. Future work could focus on refining the developed methodology by exploring more advanced regression techniques, that incorporate object segmentation and independent scale estimation for different image regions, alongside occlusion handling. This would enable more accurate depth scaling, improve metric precision, and address the complexities in the relationships among objects, as opposed to our linear regression approach.